OpenStudy (anonymous):

write phrase as an algebraic expression 35 multiplied by the quantity r less 45

OpenStudy (saifoo.khan):

\[35 \times r < 45\] \[35r < 45\] \[r < \frac{45}{35}\] \[r < \frac 97\]

OpenStudy (anonymous):

"35 multiplied by the quantity r" is: 35r and "less 45" means subtract 45 and is: -45 putting the two together we get: 35r-45